OpenStudy (anonymous):

Solve the following system using the substitution method. x + 6y = –5 2x – 9y = –17

OpenStudy (anonymous):

OK... do so :D x = -5-6y from 1st equation.... substitute to the second one -10-12y-9y=-17 21y=7 y=3 => x=-5-6*3= -23